find the perimeter of triangle ABC A(-4,-1)B(4,5)C(4,-2) I know the answer is 25.0623 by the help of one of the tutors here but when I put it in the answer block for the acellus site it says its wrong I tried using 25 first and it was wrong I then tried using the whole 25.0623 and it still said I was wrong
Answer by Alan3354(69443)

find the perimeter of triangle ABC A(-4,-1)B(4,5)C(4,-2) I know the answer is 25.0623 by the help of one of the tutors here but when I put it in the answer block for the acellus site it says its wrong I tried using 25 first and it was wrong I then tried using the whole 25.0623 and it still said I was wrong
------------
AB = sqrt(8^2 + 6^2) = 10
AC = sqrt(8^2 + 1) = sqrt(65) =~ 8.0623
BC = sqrt(0 + 7^2) = 7
Perimeter = 25.0623
==================
Does it ask for the nearest 10th? or 100th?